# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v6

@@ Abstract
We address the applicability of a hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver (CQM) as well as with CPLEX, for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments. All the requirements posed by the operator of the network included in the model. The computational results demonstrate the readiness for application and the benefits of quantum-classical hybrid solvers in a realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
@@ 3 Problem description
Our model is mainly macroscopic while it contains some microscopic features, including station technology, track occupancy, rolling stock circulation, and shunting. Concerning switching zones, as no adjusting routes are modeled, independent parts of switching zones can be considered, thus parallel entries and departures can be performed.
@@ 4.1.1 Sets
$\hat{S}$ stations
$S$ decision stations
$J$ trains
@@
Let $\hat{S}$ denote the set of all stations, and $S\subset\hat{S}$ the decision stations which have been selected in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. The sets $\mathcal{S}_{j}$ list decision stations only and they are ordered. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following.
@@
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$ (given trains’ paths fixed, $s^{\prime}$ is uniquely defined by $j^{\prime}$ and $s$). For these train pairs, the respective constraints will ensure that they do not pass the same switch in an interlocking area at the same time upon departure.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same switch in an interlocking area at the same time upon arrival, which is also enforced by the interlocking system. The possibility for M-O between the current and previous decision stations (e.g. using multiple tracks) is to be modeled specifically, hence the distinction between these sets and the $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$-s.
@@
  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction, and they cannot meet and overtake between the subsequent stations.
  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs ($s$, $s^{\prime}$) that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. (Note that $s^{\prime}$ is uniquely defined by $j$, $j^{\prime}$ and $s$ because the routes of trains are fixed.)
@@ 4.2.1 Objective function
where $s^{*}$ is the last element of $S_{j}$, and $w_{j}$ are weights for each train representing its priority. The weights have to be chosen in such a way that they reflect the rules and practices of the local dispatching system. As these rules are often formulated only qualitatively, the particular values are somewhat arbitrary; the model can be fine-tuned in this respect. (In Section 6.1 we provide more details of the weight assignments for Polish PKP PLK.) The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as, e.g., the weighted sum of secondary delays on all stations, is straightforward.
@@ 4.2.2 Constraints
$t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s),\ \forall j\in\mathcal{J},\ \forall s\in S_{j}.$ (10)
@@
$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\ {\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}},\$ (11)
@@
$\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s};$ (12)
@@
$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}};$ (13)
@@
$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}};$ (14)
@@
$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}},$ (15)
@@
$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s),\ \forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}.$ (16)
@@
$\begin{split}y^{(\text{in})}(j,j^{\prime},s^{\prime})=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(s,s^{\prime})\in S\times S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{\text{in}},\text{MO})}_{s^{\prime},s}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}.\end{split}$ (18)
@@ 4.2.3 Scaling analysis
Now let us estimate the number of constraints. Considering the headway constraints in Equation (4), there are two such constraints for each independent $y^{(\text{out})}$ variable. (Because of the symmetry property in Equation (19), only half of the $y^{(\text{out})}$ variables are independent).
Along the same arguments, in case of the constraints in Equation (11), i.e. station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each independent $z$ variable. It is more complicated to estimate the number of interlocking area occupation constraints in Equations (4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables. A good assumption is to consider $2$ constraints for each independent $y^{(\text{out})}$ variable, $2$ constraints for each independent $y^{(\text{in})}$ variable, and $2$ constraints for each independent $z$ variable. Further, we expect one minimal dwell time constraint per train and station, resulting in #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case. Concerning train order constraints in Equations (17) and (18), there can be one constraint per train and station, which yields additional #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case. As we do not consider splitting and joining trains on the route, at most one rolling stock circulation constraint arises per train, yielding at most #$\#\mathcal{J}$ such constraints. In summary, the total number of constraints obeys the following (worst-case) upper bound:
@@
#$\#constr.\leq 6\#\mathcal{J}^{2}\#\mathcal{S}-4\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (24)

### v7

@@ Abstract
We address the applicability of a hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver (CQM) as well as with CPLEX, for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments. All the requirements posed by the operator of the network are included in the model. The computational results demonstrate the readiness for application and the benefits of quantum-classical hybrid solvers in a realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
@@ 4.1.1 Sets
$\hat{\mathcal{S}}$ stations
$\mathcal{S}$ decision stations
$\mathcal{J}$ trains
@@
Let $\hat{\mathcal{S}}$ denote the set of all stations, and $\mathcal{S}\subset\hat{\mathcal{S}}$ the decision stations which have been selected in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. The sets $\mathcal{S}_{j}$ list decision stations only and they are ordered. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following.
@@
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$ (given trains’ paths fixed, $s^{\prime}$ is uniquely defined by $j^{\prime}$ and $s$). For these train pairs, the respective constraints will ensure that they do not pass the same switch in an interlocking area at the same time upon departure.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same switch in an interlocking area at the same time upon arrival, which is also enforced by the interlocking system. The possibility for M-O between the current and previous decision stations (e.g. using multiple tracks) is to be modeled specifically, hence the distinction between these sets and the $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$-s.
@@
  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}\times\mathcal{J}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction, and they cannot meet and overtake between the subsequent stations.
  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}\times\mathcal{J}\right)$ are those of all station pairs ($s$, $s^{\prime}$) that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. (Note that $s^{\prime}$ is uniquely defined by $j$, $j^{\prime}$ and $s$ because the routes of trains are fixed.)
@@ 4.2.1 Objective function
where $s^{*}$ is the last element of $\mathcal{S}_{j}$, and $w_{j}$ are weights for each train representing its priority. The weights have to be chosen in such a way that they reflect the rules and practices of the local dispatching system. As these rules are often formulated only qualitatively, the particular values are somewhat arbitrary; the model can be fine-tuned in this respect. (In Section 6.1 we provide more details of the weight assignments for Polish PKP PLK.) The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as, e.g., the weighted sum of secondary delays on all stations, is straightforward.
@@ 4.2.2 Constraints
$t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s),\ \forall j\in\mathcal{J},\ \forall s\in\mathcal{S}_{j}.$ (10)
@@
$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\ {\forall s\in\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}},\$ (11)
@@
$\forall s\in\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s};$ (12)
@@
$\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}};$ (13)
@@
$\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}};$ (14)
@@
$\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}},$ (15)
@@
$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s),\ \forall s\in\mathcal{S}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}.$ (16)
@@
$\begin{split}y^{(\text{in})}(j,j^{\prime},s^{\prime})=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(s,s^{\prime})\in\mathcal{S}\times S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{\text{in}},\text{MO})}_{s^{\prime},s}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}.\end{split}$ (18)
@@ 4.2.3 Scaling analysis
Now let us estimate the number of constraints. Considering the headway constraints in Equation (4), there are at most two such constraints for each independent $y^{(\text{out})}$ variable. (Because of the symmetry property in Equation (19), only half of the $y^{(\text{out})}$ variables are independent).
Along the same arguments, in case of the constraints in Equation (11), i.e. station track occupancies, there are at most $2$ of these for each independent $y^{(\text{out})}$ variable. Analogously, for the single-line track occupancy constraints in Equation (9) there are at most $2$ constraints per each independent $z$ variable. Furthermore, in the worst case there are: $2$ constraints for each independent $y^{(\text{out})}$ variable from Equation (4.2.2), $2$ constraints for each independent $y^{(\text{out})}$ variable from Equation (4.2.2), $2$ constraints for each independent $z$ variable from Equation (4.2.2), and $2$ constraints for each independent $y^{(\text{in})}$ variable from Equation (4.2.2). Further, we expect one minimal dwell time constraint per train and station, resulting in #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case. Concerning train order constraints in Equations (17) and (18), there can be one constraint per train and station, which yields additional #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case. As we do not consider splitting and joining trains on the route, at most one rolling stock circulation constraint arises per train, yielding at most #$\#\mathcal{J}$ such constraints. In summary, the total number of constraints obeys the following (worst-case) upper bound:
@@
#$\#constr.\leq 7\#\mathcal{J}^{2}\#\mathcal{S}-5\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (24)